OpenStudy (anonymous):

Simplify the following expression. 3x - 6y + 2z - 5x - 3y + 8z -2x + 9y + 10z -2x + 9y - 10z -2x - 9y + 10z 2x - 9y + 10z

OpenStudy (anonymous):

well first thing you want to do is combine like terms it pretty much gives you the combined terms already so now all you have to do is bring down the signs to solve.

OpenStudy (anonymous):

I'll help you out by narrowing down the answers it isn't the last 2

OpenStudy (anonymous):

show that the iterated integral \[\int\limits_{0}^{1\int\limits_{\int\limits_{0}^{?}}^{n-1}}...\int\limits_{0}^{2\left( x1+x2+...+xn \right)}\] =n!n(n+1)/4